OpenStudy (anonymous):

Rationalize the numerator of: (((x)^(1/2))-((x+h)^(1/2)))/(h((x)^(1/2))((x+h)^(1/2)) I got: 1/ (((x^2+xh)^(1/2))+((x^3+2x^2h+xh^2)^(1/2))) But am really unsure about the denominator...

OpenStudy (anonymous):

my eyes!!! too many parenthesisesess.

OpenStudy (anonymous):

\[(\sqrt{x}-\sqrt{x+h})/(h \sqrt{x}\sqrt{x+h})\]

OpenStudy (anonymous):

i'm kidding... dont need to...

OpenStudy (anonymous):

Sorry am clearing it up now that is the question...

OpenStudy (anonymous):

Oh okay, lol... is that the same answer that you got?

OpenStudy (anonymous):

\[1/ (\sqrt{x^2+xh}+\sqrt{x^3+2x^2h+xh^2}) \]

OpenStudy (anonymous):

just to clear it up.. \[\large \frac {\sqrt{x}-\sqrt{x+h}}{h \sqrt {x} \sqrt{x+h}}\]

OpenStudy (anonymous):

shouldn't that be a -1 at the top?

OpenStudy (anonymous):

yes in the answer...

OpenStudy (anonymous):

still doing it... hang on..

OpenStudy (anonymous):

okay :)

OpenStudy (dumbcow):

yeah you should get \[\frac{-1}{\sqrt{x}\sqrt{x+h}(\sqrt{x}+\sqrt{x+h})} = \frac{-1}{x \sqrt{x+h} + (x+h)\sqrt{x}}\]

OpenStudy (anonymous):

i left it in factored form (the denominator): h*sqrt(x(x+h)) * (sqrt(x+h)+sqrt(x))

OpenStudy (anonymous):

talk about parenthesisises.

OpenStudy (anonymous):

how'd you get rid of the h if you left it factored

OpenStudy (anonymous):

is it okay for me to expand it that way?